# Track4World: Feedforward World-centric Dense 3D Tracking of All Pixels

### 03 Mar 2026 v1

@@ 2.1 Video Geometry Estimation
Video geometry estimation aims to generate temporally consistent point maps from videos. Early works focus on optimization-based refinement, e.g., RobustCVD [kopf2021robust] and CasualSAM [zhang2022structure], which optimize depth and camera parameters using geometric constraints. Recent differentiable SLAM methods, such as MegaSaM [li2024megasam], Uni4D [yao2025uni4d], and ViPE [huang2025vipe], incorporate depth priors and correspondence maps for video-level point map prediction. A second line of research explores data-driven approaches [wang2024dust3r, lu2024align3r, zhang2024monst3r, chen2025easi3r, leroy2024grounding, yang2025fast3r, wang2025vggt, jiang2025geo4d, wang2025continuous, xiong2025human3r, chen2025ttt3r, zhuo2025streaming, li2026unish], leveraging world-centric geometry for robust and consistent estimation. However, VGGT [wang2025vggt] demonstrates that camera-centric depth and pose often outperform world-centric representations. Consequently, recent methods, including GeometryCrafter [xu2025geometrycrafter], Pi3 [wang2025pi], and MapAnything [keetha2025mapanything], focus on camera-centric predictions. Our approach aligns with this trend, adopting a camera-centric representation while focusing on joint geometry and motion estimation.
@@ 3.1 Overview
Given a video sequence consisting of $T$ frames, $\{\mathbf{I}_{i}\in\mathbb{R}^{H\times W\times 3}\mid i=1,\ldots,T\}$, Track4World first predicts dense 3D scene flow between arbitrary pairs of source and target timesteps $(i,j)$. It then fuses these pairwise 3D scene flows to construct a holistic 3D track for every pixel within the world-centric coordinate system. As illustrated in Fig. 2, rather than treating 3D tracking as an isolated spatial problem, we initially extract global scene representations, including geometric features, camera-centric point clouds, and camera poses, using a finetuned geometry encoder initialized from recent state-of-the-art feedforward 3D reconstruction models [wang2025pi, lin2025depth, wang2025moge1], like Pi3 or Depth Anything v3 (DA3), (see supplementary material for details). The core of our framework is a scene flow decoder that builds upon global geometric representations to estimate scene flow. This decoder operates in a sparse-to-dense manner to reduce computational overhead, which introduces a novel 2D-to-3D correlation module that entirely avoids expensive 3D spatial correlation, enabling a 2D-3D joint supervision strategy to circumvent the severe scarcity of 3D ground-truth annotations.
@@ 4.6 Ablation Study
Efficiency of motion decoder. Tab. 8 compares the inference time (Time), memory consumption (Mem.), and parameter count (Parm.) on 16-frame sequences from the ADT [pan2023aria] dataset. First, pairwise scene flow approaches, such as POMATO [zhang2025pomato] and ZeroMSF [liang2025zero], natively support dense tracking by computing flow independently for every pixel. However, unlike our efficient sparse-to-dense formulation, their per-pixel prediction approach incurs significantly higher computational cost, resulting in increased latency without providing a clear advantage in memory efficiency. On the other hand, STV2 [xiao2025spatialtrackerv2] relies on a traditional 3D spatial correlation mechanism for iterative trajectory estimation. Its high computational complexity and large parameter count impose a severe bottleneck. Consequently, STV2 is restricted to sparse tracking and suffers from Out-of-Memory (OOM) errors when extended to dense tracking scenarios. To explicitly validate that this bottleneck lies in the traditional 3D computation, we replace our proposed 2D-to-3D correlation module with the aforementioned traditional 3D mechanism (Ours w/o 2D-to-3D). As shown in the gray rows of Tab. 8, this variant similarly crashes with an OOM error under dense tracking settings. Overall, by bypassing the prohibitive cost of explicit 3D correlations, our method achieves superior efficiency in runtime, memory footprint, and model size, all while natively supporting dense tracking and maintaining state-of-the-art performance.
@@ 4.7 Qualitative Visualization
To demonstrate our method’s effectiveness, Fig. 4 visualizes the tracking and flow results across five aspects: (1) 1st-frame 2D tracking, where consistent colors denote temporal pixel correspondences. (2) Scene flow, contrasting the original RGB point cloud with the flow-transformed geometry (rainbow). (3) 1st-frame 3D tracking, featuring uniformly downsampled camera-centric trajectories for rendering efficiency. (4) All-frame 3D tracking, capturing dense camera-centric trajectories of both existing and newly emerging objects. (5) World-centric tracking, mapping trajectories to a global coordinate system. This effectively decouples camera ego-motion from object dynamics, yielding spatially stable backgrounds and physically coherent absolute motions for dynamic entities. For dynamic visualizations, please refer to the supplementary video.
Figure 4: Qualitative results on diverse in-the-wild videos.
@@ Appendix I More Visualization
To further demonstrate the capability of our framework, we provide additional qualitative results, including scene flow estimation, dense first-frame 2D tracking, dense first-frame 3D tracking, per-pixel tracking across all frames, and world-centric dense tracking across all frames, as shown in Fig.S2, Fig.S4, Fig.S5, Fig.S6 and Fig.S7. (1) Scene Flow Visualization. We first present pairwise scene flow predictions between two frames. Scene flow characterizes the 3D motion of each point and serves as an essential cue for understanding dynamic geometries. In our visualization, the deformed point maps (colored  ) highlight the predicted per-point displacements. Compared with the current state-of-the-art method ZeroMSF [liang2025zero], our results exhibit smoother spatial transitions and significantly higher temporal consistency in both geometry and motion. This demonstrates that our deformation field models coherent 3D dynamics more effectively, even in complex and rapidly changing scenes. (2) Dense 2D Tracking from the First Frame. The second group shows dense 2D correspondences anchored at the first frame. Pixels with the same color trace the evolution of the same point across time, revealing long-term and fine-grained motion patterns throughout the sequence. (3) Dense 3D Trajectories from the First Frame. We further visualize reconstructed 3D trajectories originating from the first frame. For clarity and rendering efficiency, we uniformly subsample the trajectories while preserving their overall motion structures and temporal smoothness. (4) Dense Per-Pixel Trajectories Across All Frames. We present dense per-pixel trajectories over the full sequence (second row), which consistently track motions even when new objects enter the scene, highlighting the robustness of our method in handling complex dynamic scenarios. (5) Dense Trajectories in the World Coordinate System. Finally, we visualize dense trajectories across all frames after transforming them into a unified global world coordinate system. This world-centric representation effectively decouples camera ego-motion from object dynamics: static background elements remain spatially stable over time, while dynamic objects exhibit coherent and physically meaningful absolute motion trajectories in 3D space. For more intuitive demonstrations and continuous motion effects, we refer the reader to the supplementary video.

### 05 Mar 2026 v2

@@ 2.1 Video Geometry Estimation
Video geometry estimation aims to generate temporally consistent point maps from videos. Early works focus on optimization-based refinement, e.g., RobustCVD [kopf2021robust] and CasualSAM [zhang2022structure], which optimize depth and camera parameters using geometric constraints. Recent differentiable SLAM methods, such as MegaSaM [li2024megasam], Uni4D [yao2025uni4d], and ViPE [huang2025vipe], incorporate depth priors and correspondence maps for video-level point map prediction. A second line of research explores data-driven approaches [wang2024dust3r, lu2024align3r, zhang2024monst3r, chen2025easi3r, leroy2024grounding, yang2025fast3r, wang2025vggt, jiang2025geo4d, wang2025continuous, xiong2025human3r, chen2025ttt3r, zhuo2025streaming, li2026unish], leveraging world-centric geometry for robust, consistent estimation. However, VGGT [wang2025vggt] shows that camera-centric depth and pose often outperform world-centric representations. Consequently, recent methods, including GeometryCrafter [xu2025geometrycrafter], Pi3 [wang2025pi], and MapAnything [keetha2025mapanything], focus on camera-centric predictions. Our approach follows this trend by employing a camera-centric representation (camera-centric point clouds and camera poses), from which world-centric reconstruction can be directly derived, while jointly estimating geometry and motion.
@@ 3.1 Overview
Given a video sequence of $T$ frames, $\{\mathbf{I}_{i}\in\mathbb{R}^{H\times W\times 3}\mid i=1,\ldots,T\}$, Track4World aims to construct a holistic 3D track for every pixel within a world-centric coordinate system (as illustrated in Fig. 2). To achieve this, rather than treating 3D tracking as an isolated spatial problem, our framework operates through a logical pipeline. First, we extract global scene representations, including geometric features, camera-centric point clouds, and camera poses, using a finetuned geometry encoder initialized from recent state-of-the-art feedforward 3D reconstruction models [wang2025pi, lin2025depth, wang2025moge1], such as Pi3 or Depth Anything v3 (DA3), (see supplementary material for details). Second, built upon these global representations, our core scene flow decoder predicts dense 3D scene flow between arbitrary pairs of source and target timesteps $(i,j)$. To reduce computational overhead, this decoder operates in a sparse-to-dense manner. Crucially, it introduces a novel 2D-to-3D correlation module that completely avoids expensive 3D spatial correlation, which in turn enables a 2D-3D joint supervision strategy to circumvent the severe scarcity of 3D ground-truth annotations. Finally, Track4World fuses these pairwise 3D scene flows to formulate the ultimate holistic 3D tracks.
@@ 4.6 Ablation Study
Efficiency of scene flow decoder. Tab. 8 compares the inference time (Time), memory consumption (Mem.), and parameter count (Parm.) on 16-frame sequences from the ADT [pan2023aria] dataset. First, pairwise scene flow approaches, such as POMATO [zhang2025pomato] and ZeroMSF [liang2025zero], natively support dense tracking by computing flow independently for every pixel. However, unlike our efficient sparse-to-dense formulation, their per-pixel prediction approach incurs significantly higher computational cost, resulting in increased latency without providing a clear advantage in memory efficiency. On the other hand, STV2 [xiao2025spatialtrackerv2] relies on a traditional 3D spatial correlation mechanism for iterative trajectory estimation. Its high computational complexity and large parameter count impose a severe bottleneck. Consequently, STV2 is restricted to sparse tracking and suffers from Out-of-Memory (OOM) errors when extended to dense tracking scenarios. To explicitly validate that this bottleneck lies in the traditional 3D computation, we replace our proposed 2D-to-3D correlation module with the aforementioned traditional 3D mechanism (Ours w/o 2D-to-3D). As shown in the gray rows of Tab. 8, this variant similarly crashes with an OOM error under dense tracking settings. Overall, by bypassing the prohibitive cost of explicit 3D correlations, our method achieves superior efficiency in runtime, memory footprint, and model size, all while natively supporting dense tracking and maintaining state-of-the-art performance.
@@
Figure 4: Qualitative results on diverse in-the-wild videos.
@@ 4.7 Qualitative Visualization
To demonstrate our method’s effectiveness, Fig. 4 visualizes the tracking and flow results across five aspects. The first four illustrate camera-centric perspectives: (1) 1st-frame 2D tracking, where consistent colors denote temporal pixel correspondences. (2) Scene flow, contrasting the original RGB point cloud with the flow-transformed geometry (rainbow). (3) 1st-frame 3D tracking, featuring uniformly downsampled camera-centric trajectories for rendering efficiency. (4) All-frame 3D tracking, capturing dense camera-centric trajectories of both existing and newly emerging objects. Distinct from the above, (5) World-centric tracking maps these trajectories to a global coordinate system. This effectively decouples camera ego-motion from object dynamics, yielding spatially stable backgrounds and physically coherent absolute motions for dynamic entities. For dynamic visualizations, please refer to the supplementary video.
@@ Appendix I More Visualization
To further demonstrate the capability of our framework, we provide additional qualitative results. Specifically, we present camera-centric results, including scene flow estimation, dense first-frame 2D tracking, dense first-frame 3D tracking, and per-pixel tracking across all frames (Fig. S2, Fig. S4, Fig. S5, and Fig. S6). We also show world-centric dense tracking across all frames in Fig. S7. (1) Scene Flow Visualization. We first present pairwise scene flow predictions between two frames. Scene flow characterizes the 3D motion of each point and serves as an essential cue for understanding dynamic geometries. In our visualization, the deformed point maps (colored  ) highlight the predicted per-point displacements. Compared with the current state-of-the-art method ZeroMSF [liang2025zero], our results exhibit smoother spatial transitions and significantly higher temporal consistency in both geometry and motion. This demonstrates that our deformation field models coherent 3D dynamics more effectively, even in complex and rapidly changing scenes. (2) Dense 2D Tracking from the First Frame. The second group shows dense 2D correspondences anchored at the first frame. Pixels with the same color trace the evolution of the same point across time, revealing long-term and fine-grained motion patterns throughout the sequence. (3) Dense 3D Trajectories from the First Frame. We further visualize reconstructed 3D trajectories originating from the first frame. For clarity and rendering efficiency, we uniformly subsample the trajectories while preserving their overall motion structures and temporal smoothness. (4) Dense Per-Pixel Trajectories Across All Frames. We present dense per-pixel trajectories over the full sequence (second row), which consistently track motions even when new objects enter the scene, highlighting the robustness of our method in handling complex dynamic scenarios. (5) Dense Trajectories in the World Coordinate System. Finally, we visualize dense trajectories across all frames after transforming them into a unified global world coordinate system. This world-centric representation effectively decouples camera ego-motion from object dynamics: static background elements remain spatially stable over time, while dynamic objects exhibit coherent and physically meaningful absolute motion trajectories in 3D space. For more intuitive demonstrations and continuous motion effects, we refer the reader to the supplementary video.